OpenStudy (anonymous):

solve the equation 13r+10=5r-46 r=???

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

8r = -56 Divide by 8 to get r..

OpenStudy (anonymous):

Thanks you guys <3

OpenStudy (anonymous):

13r+10=5r-46 13r+10-5r=5r-46-5r 8r+10=-46 8r+10-10=-46-10 8r=-56 8r/8=-56/8 r=-7

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

13r+10=5r-46 -10 -10 13r=5r-56 -5r -5r 8r=-56 r=-7